OpenStudy (anonymous):

-√72 answer and explain

OpenStudy (anonymous):

\[72=2^3\times 3^2\] so \[\sqrt{72}=2\times 3\times \sqrt{2}\] therefore \[-\sqrt{72}=-6\sqrt{2}\]

OpenStudy (anonymous):

might be easier just to say \[-\sqrt{72}=-\sqrt{36\times 2}=-6\sqrt{2}\]